OpenStudy (anonymous):

Find the first term of the following geometric sequence. _____, 39, 117, 351, 1,053, ...

OpenStudy (anonymous):

Since it is a geometric sequence, there has to be a common ratio between the terms. Divide a term by the previous term to figure out what it is.

OpenStudy (anonymous):

So: \[\frac{117}{39} = 3\] the ratio is 3, so divide 39 by 3 to get your answer :)

OpenStudy (anonymous):

but the appex thing said the answer is 13

OpenStudy (anonymous):

yeah, (39/3) = 13 :)

OpenStudy (amistre64):

39/3 = 13 right?

OpenStudy (anonymous):

divide 117/39 =3 thus 3 is the common ratio \[T_{n}=ar ^{n-1}\]

OpenStudy (anonymous):

ok, i c

OpenStudy (anonymous):

n=1 r=3 use the above formula to look 4 a